OpenStudy (anonymous):

What is the value of 16 – 2(3 • 5 + 1)?

OpenStudy (compassionate):

PEMDAS: 16 - 2(15 + 1) 16 - 2(16) -42(16) = -672

OpenStudy (compassionate):

No, 1 16 -2(16) 16 - 32 -16

OpenStudy (anonymous):

The options are 8 1 -16 and -21

OpenStudy (ghazi):

16-(7+2)= 16-9=7

OpenStudy (ghazi):

well i thought it's 3.5 but it's 3*5 so answer is -16

OpenStudy (anonymous):

Are you sure?